find the x- and y- intercepts.
8x-2y= 16
x-intercept (x,y) = (__________,_________)
y-intercept (x,y) = (__________,_________)
Use the intercepts to graph the equation.
Please help!!

Found 2 solutions by haileytucki, rfer:
Answer by haileytucki(390)

8x-2y=16
To find the x-intercept, substitute in 0 for y and solve for x.
8x-2(0)=16
Multiply -2 by each term inside the parentheses.
8x+0=16
Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression.
8x=16
Divide each term in the equation by 8.
(8x)/(8)=(16)/(8)
Simplify the left-hand side of the equation by canceling the common factors.
x=(16)/(8)
Simplify the right-hand side of the equation by simplifying each term.
x=2
To find the y-intercept, substitute in 0 for x and solve for y.
8(0)-2y=16
Multiply 8 by each term inside the parentheses.
0-2y=16
Combine all similar expressions.
-2y=16
Divide each term in the equation by -2.
-(2y)/(-2)=(16)/(-2)
Simplify the left-hand side of the equation by canceling the common factors.
y=(16)/(-2)
Solve the equation.
y=-8
These are the x and y intercepts of the equation 8x-2y=16.
x=2, y=-8

Answer by rfer(12662)

x int (2,0)
y int (0,-8)